1. ( Multinat PLC has asked you to evaluate the following project for the production of a new product. The firm has already spent £100,000 on marketing consultant fees to estimate potential sales of the new item. You are going to charge a fee of £10,000 to undertake the project evaluation. You know that it will initially be necessary to invest E1m in a piece of new machinery. It has been estimated that expected sales are 100 items of the new product in years 1, 2 and 3 but nothing thereafter. In year 1 you expect items to sell at £10,000 per item in nominal terms. The nominal total costs of production will be £500,000 in year 1. In addition, head office costs will rise because it will be necessary to recruit one new employee on a three-year contract to oversee the project at a starting cost to the firm of £30,000. However, from an accounting perspective, the project will be allocated with £50,000 of head office costs in the first year Question 1 continued/... Page 1 of 4 You expect sales revenues and all costs and allocations to rise in line with inflation for years 2 and 3, Inflation is expected to run at 2% per annum. At the end of the third year the machinery has no resale value. Identify the relevant incremental pre-tax cash flows associated with the project (25 marks) (b) Assume that Multinat PLC is a highly profitable company. It pays corporation tax at a rate of 30%. For calculating capital allowances for taxation purposes, assume that the new piece of machinery can be depreciated at 25% on a straight-line basis with the first allowance coming immediately (year 0). Identify all the relevant post- tax cash flows associated with this project. (25 marks) (c) You are trying to estimate the cost of capital at which to discount this project. You know that the annualised risk-free rate is 2% and estimate that the annualised equity premium is 4%. You have found a suitable comparison company with an equity beta of 1.2. This comparison company has 1 million shares in issue with a current share price of £6.00 and total debt of £3,000,000. Multinat PLC can borrow at a pre-tax rate of 3%, has a target debt to debt + equity ratio (DDE) of 0.4 and is planning on funding this project entirely with retained earnings. Assuming that the debt betas of both the comparison company and Multinat PLC are zero, what is the appropriate post-tax weighted average cost of capital? (30 marks) 1. (d) What are the discounted payback period and NPV of this project? Should Multinat PLC go ahead with the project?
Answer #1
Year 0 1 2 3
CAPEX cash flows:
1.Cost of new machinery -1000000
2.Operating Cash flows
i.Sale items 100 100 100
ii. Sale value(a*10000) 1000000 1020000 1040400
iii.Less:Production costs -500000 -510000 -520200
iv.Less:Salary of new employee -30000 -30600 -31212
a. Incremental pre-tax cash flows(ii-iii-iv) 470000 479400 488988
v.Less: Incl.Tax at 30% -141000 -143820 -146696
vi.Add: Depn. Tax shield(1000000/4*30%) 75000 75000 75000 75000
vii.Incremental post-tax cash flows(a-v+vi) 75000 404000 410580 417292
b. Total Incl.post-tax cash flows(1+vii) -925000 404000 410580 417292
3.PV F at 4.92%(1/1.0492^n) 1 0.95311 0.90841 0.86582
4. PV at 4.92%(b*3) -925000 385055 372976 361297
5. NPV (sum row 4) 194329
Discounted pay-back period:
4. PV at 4.92%(b*3) -925000 385055 372976 361297
5. Cumulative p/b -925000 -539945 -166968 194329
Disc.Pay-back period= 2+(166968/361297)=
2.46
Years
d. Multinat PLC can go ahead with the project as it returns POSITIVE NPV
Also it pays back within 3 years.
WACC Calculations:
As per CAPM
Cost of Retained eranings,K(re) =RFR+(Beta*Market Risk Premium)
K(re)=2%+(1.2*4%)=
6.80%
After-tax cost of debt=3%*(1-30%)=
2.10%
So,
Post-tax WACC for Multinat PLC is:
WACC=(wd*kd)+(we*ke)
ie.(40%*2.10%)+(60%*6.80%)=
4.92%